OpenStudy (lgbasallote):

Find the domain: \[h(x) = \sqrt{4-x} + \sqrt{x^2 - 1}\]

OpenStudy (lgbasallote):

i smell a race....

Parth (parthkohli):

Hint: you can't use positive numbers more than 4.

OpenStudy (lgbasallote):

^i wish you would have let me try first....

OpenStudy (lgbasallote):

^@Jonask

OpenStudy (lgbasallote):

^and then you did it again @Jonask

OpenStudy (anonymous):

sorry,like you said its a race

OpenStudy (lgbasallote):

i hate races.....

OpenStudy (swissgirl):

Ok just try to see when each square root is negative

OpenStudy (lgbasallote):

anyway....

OpenStudy (lgbasallote):

i suppose \(\sqrt{x^2 - 1}\) will have no restrictions?

OpenStudy (swissgirl):

yes x=0

OpenStudy (lgbasallote):

hmm oh yeah

OpenStudy (lgbasallote):

didn't think of that

OpenStudy (lgbasallote):

then \(\sqrt{4-x}\) would be 4 above?

OpenStudy (swissgirl):

I would say greater than 4

OpenStudy (anonymous):

@lgbasallote was that the question of 100 level???

OpenStudy (lgbasallote):

greater than 4...4 above...same shiz

OpenStudy (anonymous):

Take Seperately and Connect it by union (U)

OpenStudy (lgbasallote):

to be the best, one needs to master the basics @sauravshakya

OpenStudy (anonymous):

\[f(x)=x^2,g(x)=\sqrt{1-x}\] domain of\[gof,fog\] domain

OpenStudy (lgbasallote):

^?

OpenStudy (anonymous):

just a question that i wont interupt you can try

OpenStudy (swissgirl):

Domain is just 1 and 0

OpenStudy (lgbasallote):

g o f would be sqrt(1 - x^2) it would only be non-negative if 0 <= x <= 1

OpenStudy (lgbasallote):

so the domain is 0<=x<=1

OpenStudy (swissgirl):

for f(g(x))

OpenStudy (lgbasallote):

f o g would be 1- x so all real numbers

OpenStudy (anonymous):

\[Domain=(-\infty,0) U (0,4]\]

Parth (parthkohli):

0 isn't included in the domain since \(\sqrt{0^2 - 1} = \sqrt{-1}\) and this is a real function.

OpenStudy (anonymous):

Lol...@ParthKohli u see it is a open Bracket.....

OpenStudy (lgbasallote):

"real" function?

Parth (parthkohli):

@Yahoo! I wasn't talking to you.

Parth (parthkohli):

Yeah, which has all ranges as real numbers.

OpenStudy (anonymous):

oh.....Sorry...) @ParthKohli

OpenStudy (lgbasallote):

now you confuse me...you're telling me they're wrong?

Parth (parthkohli):

swissgirl Best Response 1 Domain is just 1 and 0 that ^^

OpenStudy (lgbasallote):

hmm seems one of the answers here is wrong then...wonder which

Parth (parthkohli):

Okay, let me ask you a question. Why didn't you include all real numbers in the domain? @lgbasallote

OpenStudy (lgbasallote):

i did

OpenStudy (lgbasallote):

when didn't i?

Parth (parthkohli):

lgbasallote Best Response 0 i suppose sqrt(x62−1) will have no restrictions? swissgirl Best Response 1 yes x=0 lgbasallote Best Response 0 hmm oh yeah

Parth (parthkohli):

0 is included in all real numbers.

OpenStudy (lgbasallote):

since when was \(\sqrt{-1}\) real @ParthKohli ?

Parth (parthkohli):

You answered your own question @lgbasallote

OpenStudy (lgbasallote):

....what exactly are you saying?

Parth (parthkohli):

ParthKohli 0 0 isn't included in the domain since sqrt(0^2 - 1) = sqrt(-1) and this is a real function. <argument happens> ParthKohli 0 Okay, let me ask you a question. Why didn't you include all real numbers in the domain? @lgbasallote lgbasallote Best Response 0 i did lgbasallote Best Response 0 when didn't i? <some more convo> lgbasallote Best Response 0 since when was √−1 real @ParthKohli ?

OpenStudy (lgbasallote):

just state your point

Parth (parthkohli):

You are contradicting yourself. First, you say that you have included all real numbers in the domain. Then, by "since when was √−1 real @ParthKohli ?," you mean that 0 is not in the domain.

OpenStudy (lgbasallote):

by the way 1) none of us said 0 is included in the domain (we were talking about restrictions) 2) sqrt -1 is not a real function

OpenStudy (anonymous):

\[D_g=(-\infty,1),D_f=\mathbb{R}\] \[fog=\left| 1-x \right|\]\[gof=\sqrt{1-x^2}\]

OpenStudy (lgbasallote):

^?

Parth (parthkohli):

No, real function means that for all domains, the range will be real. We have to get a domain which keeps all ranges real.

OpenStudy (lgbasallote):

and your point is?

Parth (parthkohli):

That 0 is not in the domain.

OpenStudy (lgbasallote):

and isn't that what we said?

Parth (parthkohli):

You said that you included all real numbers in the domain. -.-

OpenStudy (lgbasallote):

when???

OpenStudy (lgbasallote):

we were talking about RESTRICTIONS

Parth (parthkohli):

ParthKohli 0 Okay, let me ask you a question. Why didn't you include all real numbers in the domain? @lgbasallote lgbasallote i did 10 minutes ago lgbasallote Best Response 0 when didn't i?

Parth (parthkohli):

I was talking about the domain then.

OpenStudy (anonymous):

\[D_{gof}=[-1,1]\] \[Dfog=x \in (-\infty,1]\]

OpenStudy (lgbasallote):

you really tend to make things big when you misread, don;t you @ParthKohli

OpenStudy (anonymous):

guys i think we can close the question i did not mean to create a big issue

OpenStudy (lgbasallote):

it has been closed a long time ago @jonask

OpenStudy (anonymous):

oh i dint notice,cos commentsn are still running